# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v3

@@ Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, while they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we reveal that CLIP’s vulnerabilities primarily stem from its tendency to encode features beyond in-dataset predictive patterns, compromising its visual feature resistivity to input perturbations. This makes its encoded features highly susceptible to being reshaped by backdoor triggers. To address this challenge, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs deep visual prompt tuning with a specially designed feature-repelling loss. Specifically, RVPT adversarially repels the encoded features from deeper layers while optimizing the standard cross-entropy loss, ensuring that only predictive features in downstream tasks are encoded, thereby enhancing CLIP’s visual feature resistivity against input perturbations and mitigating its susceptibility to backdoor attacks. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters in CLIP, yet it significantly outperforms state-of-the-art defense methods, reducing the attack success rate from 89.70% to 2.76% against the most advanced multimodal attacks on ImageNet and effectively generalizes its defensive capabilities across multiple datasets.
@@ 1 Introduction
Figure 1: Perturbation Resistivity (PR) across different layers of the encoders under various perturbations, including the trigger pattern. A higher PR value indicates less sensitivity to input perturbation. Four encoder variants are evaluated on ImageNet: (a) ViT in backdoored CLIP radford2021clip . (b) ViT in backdoored CLIP tuned with Visual Prompt Tuning (VPT) on ImageNet. (c) ViT of the backdoored CLIP tuned with our RVPT on ImageNet. (d) ViT trained exclusively on clean ImageNet paszke2019pytorch . Specifically, CLIP is backdoored by the trigger of BadCLIP liang2024badclip . Detailed experimental settings and additional PR results for CLIP backdoored by other triggers can be found in Appendix A.
@@
CLIP’s vulnerabilities to backdoor attacks have inspired the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters bansal2023cleanclip or require the availability of poisoned data yang2023safeclip ; ishmam2024semantic ; yang2023robust , which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) jia2022visual , utilizing only a small portion of downstream clean samples to encourage the model to ignore the trigger feature in the poisoned images. However, simply adopting VPT proves ineffective experimentally (see Figure 1(b)), as the trigger feature cannot be extracted from clean samples, preventing VPT from learning how to suppress them.
@@
where $f^{v}(\cdot)$ denotes the visual feature extractor. A lower PR value implies larger visual embedding drift due to input perturbations, hence greater sensitivity and higher susceptibility to backdoor triggers. To illustrate this vulnerability of CLIP, we apply a suite of perturbations to ImageNet samples and plot the layer-wise mean PR for several model variants. The results in Figure 1 show that the deeper layers of standard CLIP (Figure 1(a)) and standard CLIP tuned with VPT (Figure 1(b)) exhibit significantly lower PR than a ViT trained exclusively on clean ImageNet (Figure 1(d)), confirming that CLIP’s embeddings are easily distorted. Because a backdoor trigger acts as a structured perturbation, this low PR allows the trigger to maliciously reshape the encoded visual features layer by layer, enabling a successful attack. We therefore regard PR as a proxy for backdoor robustness, which means the backdoor-robust CLIP should have a higher PR than backdoor-vulnerable CLIP (i.e., the standard CLIP). Our analysis suggests that CLIP’s low PR originates from its large-scale, noisy pre-training dataset: the encoder captures many off-dataset, non-predictive features, which magnify the impact of perturbations. In contrast, the ImageNet-supervised ViT only learns in-dataset predictive features of ImageNet, prohibiting it from encoding irrelevant features beyond ImageNet and achieving a much higher PR. Hence, an effective defense (ours shown in Figure 1(c)) should suppress the irrelevant features CLIP inherently captures and guide it to focus on in-dataset predictive patterns, thereby enhancing both its perturbation resistivity and backdoor robustness.
Inspired by this finding, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), which improves CLIP’s PR by making it exclusively encode the predictive features in the downstream tasks in a few-shot manner. As illustrated in Figure 2, RVPT jointly optimizes two objectives: (1) a multi-layer feature-repelling loss that minimizes the cosine similarity between prompted features and original CLIP features at each layer, and (2) cross-entropy loss for downstream task accuracy. This mechanism forces the model to only extract features that contribute to CE loss optimization during training, while ignoring non-essential features (e.g., noise or trigger pattern). As a result, Figure 1(c) shows that RVPT significantly improves the PR of CLIP on ImageNet, hence effectively defending against the state-of-the-art multimodal backdoor attack. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
@@ 3 The Proposed Approach
Overview. In this section, we first outline the basic settings of the threat model and defender’s goal in Section 3.1, followed by a detailed illustration of how the adversary launches backdoor attacks on CLIP in Section 3.2. Eventually, in Section 3.3, we present Repulsive Visual Prompt Tuning, designed to defend backdoored CLIP against multiple multimodal backdoor attacks.
@@ 3.2 Backdoor Attacks on CLIP
Upon training with $\widetilde{\mathcal{D}}$, the encoders of CLIP will be backdoored, denoted as $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$. For the backdoored version of CLIP, the trigger pattern $\Theta$ will have a spurious correlation with the proxy caption of the target class $y_{t}$. Thus, during the inference stage, when the model encounters the images containing the trigger pattern, its prediction on the target class $y_{t}$ will increase significantly. In the meantime, the backdoored CLIP performs normally on benign inputs.
@@ 3.3 Repulsive Visual Prompt Tuning
Formally, assume the defender is given the backdoored CLIP, denoted by $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$ alongside a small set of clean labeled samples $\{\boldsymbol{x}_{i},y_{i}\}^{N}_{i=1}$, $y_{i}\in\mathcal{Y}=\{1,\dots,C\}$. In particular, our focus is on CLIP with the Vision Transformer dosovitskiy2020vit as its visual encoder. For clean input $\boldsymbol{x}_{i}$, the feature representation at the $d$-th layer of CLIP’s visual encoder is $\boldsymbol{f}^{d}_{i}=[\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i}]$, where $\boldsymbol{c}^{d}_{i}\in\mathbb{R}^{1\times d_{v}}$ denotes the class embedding that will finally be projected to the vision-language joint space and $\boldsymbol{e}^{d}_{i}\in\mathbb{R}^{M\times d_{v}}$ are the patch embeddings. Specifically, $M$ and $d_{v}$ are hyperparameters representing the number of patches and dimensions of the visual embedding.
Assume the visual encoder has $L$ layers, then starting from the $D$-th layer, we introduce $L-D$ repulsive visual prompts, each of length $b$, defined as $\{\boldsymbol{p}^{d}\in\mathbb{R}^{b\times d_{v}}\}_{d=D}^{L-1}$. These prompts are concatenated with the features from each subsequent block recursively, following the formulation:
@@ 4.1 Experiment Settings
Models and benchmarks. Following the settings of bansal2023cleanclip , to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M radford2021clip using 500K image-text pairs from the CC3M dataset cc3m , out of 1500 are poisoned with various types of backdoor attacks: BadNet gu2019badnets , Blended chen2017blend , ISSBA issba , WaNet nguyen2021wanet , TrojVQA walmer2022dual and BadCLIP liang2024badclip . We utilize ImageNet deng2009imagenet to evaluate the performance of defense methods against all the aforementioned attacks, with the target class of the attacks set to banana. Additionally, we use Caltech101 fei2004learning and OxfordPet parkhi2012cats (a fine-grained dataset) to evaluate defense methods against BadNet, Blended, and WaNet. These attacks target “accordion” in Caltech101 and “samoyed” in OxfordPet, respectively. We compare RVPT with Linear Probe and CleanCLIP bansal2023cleanclip . The implementation details of the backdoor attack methods, backdoor defense methods, and benchmarks will be illustrated in Appendix B.
@@
Implementation details. Unless otherwise specified, we will use ViT/B32 dosovitskiy2020vit as CLIP’s visual encoder. The prompt is initialized from a zero mean Gaussian distribution. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP radford2021clip : (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule. We conduct experiments on eight NVIDIA RTX 3090 GPUs and the computational expenses are shown in Appendix C.
@@ 4.2 The Effectiveness of RVPT
[…]
@@
[…]
@@
In Table 1 and 2, we compare RVPT with CleanCLIP and Linear Probe against six attacks on ImageNet and three attacks on Caltech101 and OxfordPets. Before poisoning, CLIP’s CA on ImageNet, Caltech101, and OxfordPets is 62.01%, 91.46%, and 86.42%, respectively.
@@
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance and requires more computing resources. Moreover, Linear Probe performs well against the first four attacks. However, it does not eliminate the trigger feature and thus performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since Linear Probe inevitably excludes CLIP’s text encoder, it cannot be used to defend against TrojVQA, which targets the multimodal task of visual question answering.
[…]
@@
Finally, RVPT effectively defends against all types of attacks while maintaining clean accuracy. It demonstrates consistently strong defensive performance across all attacks and, notably, significantly reduces the ASR of BadCLIP liang2024badclip , which is hard to defend by current defense methods. Moreover, as shown in Table 3, RVPT’s classification accuracy on poisoned images is also much higher than other defense methods, highlighting its enhanced capability in maintaining correct predictions under backdoor attacks. Moreover, we observed that even a non-backdoored CLIP experiences a significant drop in PA, further supporting our claim of its low visual feature resistivity to input perturbation.
[…]
@@
[…]
@@
[…]
@@ 4.3 The Generalization of RVPT’s Defensive Ability
Performance on the emerging target class. We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its defense capability for unseen target classes in the dataset, and the results are shown in Table 4. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table 4 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully filtered out the distinctive out-of-dataset pattern. Moreover, we can see that CA maintains except OxfordPets, which decreases by around 2%. The CA drop is due to this dataset containing only 37 classes, so the absence of 3% of the data can impact its accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
Performance on cross-dataset transfer. To evaluate RVPT’s defensive ability across datasets, we tuned CLIP with RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table 5. For all three attacks, although ASR increases slightly, it remains at a very low level. Additionally, RVPT’s accuracy decreases and occasionally falls below that of CleanCLIP. We conjecture the reason is that ImageNet does not share some of the in-dataset predictive features with Caltech101 or OxfordPets. In some cases, The non-essential feature unlearned in ImageNet may be the discriminative features in them and some of their non-essential features are not removed during tuning in ImageNet. Thus, there is a slight CA loss and ASR increase after RVPT. Nevertheless, the in-dataset predictive features of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset also contributes to the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still outperforms CleanCLIP.
Performance on cross-domain transfer. To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 imegenetv2 , ImageNet-A imageneta , ImageNet-R imagenetr , and ImageNet-S imagenets . The result is shown in Table 6. RVPT can generalize its defensive ability across different domains. We also observe that RVPT will experience a loss of recognition and defense performance with domain shifts. In particular, it shows a significant CA drop on ImageNet-A, which reflects its effectiveness in focusing on in-dataset predictive features and ignoring other features, since ImageNet-A intentionally collects natural adversarial samples for ImageNet classifiers. Notably, unlike adversarial training bai2021recent , RVPT is not designed to defend against adversarial samples in ImageNet-A. The detailed comparison between RVPT and adversarial training is in Appendix E.

### v4

@@ Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, while they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we reveal that CLIP’s vulnerabilities primarily stem from its tendency to encode features beyond in-dataset predictive patterns, compromising its visual feature resistivity to input perturbations. This makes its encoded features highly susceptible to being reshaped by backdoor triggers. To address this challenge, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs deep visual prompt tuning with a specially designed feature-repelling loss. Specifically, RVPT adversarially repels the encoded features from deeper layers while optimizing the standard cross-entropy loss, ensuring that only predictive features in downstream tasks are encoded, thereby enhancing CLIP’s visual feature resistivity against input perturbations and mitigating its susceptibility to backdoor attacks. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters in CLIP, yet it significantly outperforms state-of-the-art defense methods, reducing the attack success rate from 89.70% to 2.76% against the most advanced multimodal attacks on ImageNet and effectively generalizes its defensive capabilities across multiple datasets. The code is publicly available in our GitHub repository: https://github.com/zhangzf01/RVPT.
@@ 1 Introduction
Figure 1: Perturbation Resistivity (PR) across different layers of the encoders under various perturbations, including the trigger pattern. A higher PR value indicates less sensitivity to input perturbation. Four encoder variants are evaluated on ImageNet: (a) ViT in backdoored CLIP radford2021clip . (b) ViT in backdoored CLIP tuned with Visual Prompt Tuning (VPT) on ImageNet. (c) ViT of the backdoored CLIP tuned with our RVPT on ImageNet. (d) ViT trained exclusively on clean ImageNet paszke2019pytorch . Specifically, CLIP is backdoored by the trigger of BadCLIP liang2024badclip . Detailed experimental settings and additional PR results for CLIP backdoored by other triggers can be found in Appendix˜B.
@@
CLIP’s vulnerabilities to backdoor attacks have inspired the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters bansal2023cleanclip or require the availability of poisoned data yang2023safeclip ; ishmam2024semantic ; yang2023robust , which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) jia2022visual , utilizing only a small portion of downstream clean samples to encourage the model to ignore the trigger feature in the poisoned images.  [^1] We include the detailed reasons for using VPT to defend backdoored CLIP in Appendix A. However, simply adopting VPT proves ineffective experimentally (see Figure˜1(b)), as the trigger feature cannot be extracted from clean samples, preventing VPT from learning how to suppress them.
@@
where $f^{v}(\cdot)$ denotes the visual feature extractor. A lower PR value implies larger visual embedding drift due to input perturbations, hence greater sensitivity and higher susceptibility to backdoor triggers. To illustrate this vulnerability of CLIP, we apply a suite of perturbations to ImageNet samples and plot the layer‑wise mean PR for several model variants. The results in Figure˜1 show that the deeper layers of standard CLIP (Figure˜1(a)) and standard CLIP tuned with VPT (Figure˜1(b)) exhibit significantly lower PR than a ViT trained exclusively on clean ImageNet (Figure˜1(d)), confirming that CLIP’s embeddings are easily distorted. Because a backdoor trigger acts as a structured perturbation, this low PR allows the trigger to maliciously reshape the encoded visual features layer by layer, enabling a successful attack. We therefore regard PR as a proxy for backdoor robustness, which means the backdoor-robust CLIP should have a higher PR than backdoor-vulnerable CLIP (i.e., the standard CLIP). Our analysis suggests that CLIP’s low PR originates from its large‑scale, noisy pre‑training dataset: the encoder captures many off‑dataset, non‑predictive features, which magnify the impact of perturbations. In contrast, the ImageNet‑supervised ViT only learns in-dataset predictive features of ImageNet, prohibiting it from encoding irrelevant features beyond ImageNet and achieving a much higher PR. Hence, an effective defense (ours shown in Figure˜1(c)) should suppress the irrelevant features CLIP inherently captures and guide it to focus on in-dataset predictive patterns, thereby enhancing both its perturbation resistivity and backdoor robustness.
Inspired by this finding, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), which improves CLIP’s PR by making it exclusively encode the predictive features in the downstream tasks in a few-shot manner. As illustrated in Figure˜2, RVPT jointly optimizes two objectives: (1) a multi-layer feature-repelling loss that minimizes the cosine similarity between prompted features and original CLIP features at each layer, and (2) cross-entropy loss for downstream task accuracy. This mechanism forces the model to only extract features that contribute to CE loss optimization during training, while ignoring non-essential features (e.g., noise or trigger pattern). As a result, Figure˜1(c) shows that RVPT significantly improves the PR of CLIP on ImageNet, hence effectively defending against the state-of-the-art multimodal backdoor attack. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
@@ 3 The Proposed Approach
Overview. In this section, we first outline the basic settings of the threat model and defender’s goal in Section˜3.1, followed by a detailed illustration of how the adversary launches backdoor attacks on CLIP in Section˜3.2. Eventually, in Section˜3.3, we present Repulsive Visual Prompt Tuning, designed to defend backdoored CLIP against multiple multimodal backdoor attacks.
@@ 3.1 Basic Settings
Clarification and practical motivation.
Although the above setting assumes the defender lacks access to the original pre-training data, this assumption is both realistic and practically important. First, the backdoor robustness of many multimodal models depends on CLIP, whose pre-training data is entirely unavailable to the public or defenders. Even though some models partially disclose data sources (e.g., LAION, CC3M), most real-world multimodal models are trained on large-scale, proprietary, or mixed Internet data, making it infeasible for defenders to reconstruct the same distribution. Recent studies also show that backdoored CLIP models can transfer their vulnerabilities to other multimodal models where CLIP serves as the visual encoder, highlighting the need to defend data-closed models. Second, even when pre-training data is partially available, leveraging it for defense remains highly challenging. Because multimodal datasets are vast and Internet-sourced, locating and removing poisoned samples is computationally prohibitive. Using a surrogate dataset for fine-tuning (e.g., CC3M) is an alternative, but it causes feature mismatches and significantly degrades defense performance, as verified experimentally. In contrast, our setting targets defenders with only a small amount of trusted downstream data—a realistic and widely encountered scenario. Such a data-limited but model-accessible defense setting frequently appears in practice. For example, enterprises or institutions may adapt pre-trained CLIP models to downstream tasks (e.g., medical imaging, rare-disease recognition, or industrial inspection) but only have few-shot data due to privacy or scarcity. In these cases, defenders can access the backdoored model but not its large-scale training data. Our RVPT method is designed for this setting: it provides an end-to-end defense that adapts CLIP to downstream tasks while mitigating backdoors using only a few clean samples.
@@ 3.2 Backdoor Attacks on CLIP
[…]
@@ 3.3 Repulsive Visual Prompt Tuning
[…]
Assume the visual encoder has $L$ layers, then starting from the $D$-th layer, we introduce LDL-D repulsive visual prompts, each of length $b$, defined as $\{\boldsymbol{p}^{d}\in\mathbb{R}^{b\times d_{v}}\}_{d=D}^{L-1}$. These prompts are concatenated with the features from each subsequent block recursively, following the formulation:
@@ 4.1 Experiment Settings
Models and benchmarks. Following the settings of bansal2023cleanclip , to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M radford2021clip using 500K image-text pairs from the CC3M dataset cc3m , out of 1500 are poisoned with various types of backdoor attacks: BadNet gu2019badnets , Blended chen2017blend , ISSBA issba , WaNet nguyen2021wanet , TrojVQA walmer2022dual and BadCLIP liang2024badclip . We utilize ImageNet deng2009imagenet to evaluate the performance of defense methods against all the aforementioned attacks, with the target class of the attacks set to banana. Additionally, we use Caltech101 fei2004learning and OxfordPet parkhi2012cats (a fine-grained dataset) to evaluate defense methods against BadNet, Blended, and WaNet. These attacks target “accordion” in Caltech101 and “samoyed” in OxfordPet, respectively. We compare RVPT with Linear Probe and CleanCLIP bansal2023cleanclip . The implementation details of the backdoor attack methods, backdoor defense methods, and benchmarks will be illustrated in Appendix˜C.
@@
Implementation details. Unless otherwise specified, we will use ViT/B32 dosovitskiy2020vit as CLIP’s visual encoder. The prompt is initialized from a zero mean Gaussian distribution. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP radford2021clip : (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule. We conduct experiments on eight NVIDIA RTX 3090 GPUs and the computational expenses are shown in Appendix˜D.
@@ 4.2 The Effectiveness of RVPT
Table 1: Main results. We report ASR ($\downarrow\%$), with CA ($\%$) shown in parentheses on ImageNet1K.
@@
Table 2: Main results. We report ASR ($\downarrow\%$), with CA ($\%$) in parentheses on Caltech101, OxfordPets.
@@
In Table˜1 and 2, we compare RVPT with CleanCLIP and Linear Probe against six attacks on ImageNet and three attacks on Caltech101 and OxfordPets. Before poisoning, CLIP’s CA on ImageNet, Caltech101, and OxfordPets is 62.01%, 91.46%, and 86.42%, respectively.
@@
Table 3: Poisoned accuracy ($\uparrow\%$), defined as the model’s classification accuracy on poisoned images.
@@
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance and requires more computing resources. Moreover, Linear Probe performs well against the first four attacks. However, it does not eliminate the trigger feature and thus performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since Linear Probe inevitably excludes CLIP’s text encoder, it cannot be used to defend against TrojVQA, which targets the multimodal task of visual question answering. Finally, RVPT effectively defends against all types of attacks while maintaining clean accuracy. It demonstrates consistently strong defensive performance across all attacks and, notably, significantly reduces the ASR of BadCLIP liang2024badclip , which is hard to defend by current defense methods. Moreover, as shown in Table˜3, RVPT’s classification accuracy on poisoned images is also much higher than other defense methods, highlighting its enhanced capability in maintaining correct predictions under backdoor attacks. Moreover, we observed that even a non-backdoored CLIP experiences a significant drop in PA, further supporting our claim of its low visual feature resistivity to input perturbation.
Table 4: Results on emerging target class. We report ASR ($\downarrow\%$), with CA ($\%$) shown in parentheses. (woTC) indicates RVPT is tuned in the dataset without the target class.
@@
Table 5: Results on cross-dataset generalization. We report ASR ($\downarrow\%$), with CA ($\%$) shown in parentheses. RVPT* means that the model is tuned on ImageNet1K using RVPT and tested on Caltech101 (target class: accordion) and OxfordPets (target class: samoyed).
@@
Table 6: Results on cross-domain generalization. We report ASR ($\downarrow\%$), with CA ($\%$) shown in parentheses. We tune CLIP with RVPT on ImageNet1K and conduct CleanCLIP on CC3M.
@@ 4.3 The Generalization of RVPT’s Defensive Ability
Performance on the emerging target class. We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its defense capability for unseen target classes in the dataset, and the results are shown in Table˜4. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table˜4 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully filtered out the distinctive out-of-dataset pattern. Moreover, we can see that CA maintains except OxfordPets, which decreases by around 2%. The CA drop is due to this dataset containing only 37 classes, so the absence of 3% of the data can impact its accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
Performance on cross-dataset transfer. To evaluate RVPT’s defensive ability across datasets, we tuned CLIP with RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table˜5. For all three attacks, although ASR increases slightly, it remains at a very low level. Additionally, RVPT’s accuracy decreases and occasionally falls below that of CleanCLIP. We conjecture the reason is that ImageNet does not share some of the in-dataset predictive features with Caltech101 or OxfordPets. In some cases, The non-essential feature unlearned in ImageNet may be the discriminative features in them and some of their non-essential features are not removed during tuning in ImageNet. Thus, there is a slight CA loss and ASR increase after RVPT. Nevertheless, the in-dataset predictive features of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset also contributes to the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still outperforms CleanCLIP.
Performance on cross-domain transfer. To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 imegenetv2 , ImageNet-A imageneta , ImageNet-R imagenetr , and ImageNet-S imagenets . The result is shown in Table˜6. RVPT can generalize its defensive ability across different domains. We also observe that RVPT will experience a loss of recognition and defense performance with domain shifts. In particular, it shows a significant CA drop on ImageNet-A, which reflects its effectiveness in focusing on in-dataset predictive features and ignoring other features, since ImageNet-A intentionally collects natural adversarial samples for ImageNet classifiers. Notably, unlike adversarial training bai2021recent , RVPT is not designed to defend against adversarial samples in ImageNet-A. The detailed comparison between RVPT and adversarial training is in Appendix˜F.